OpenStudy (anonymous):

SOLVE IT PLEASE ‎2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=??

OpenStudy (anonymous):

pls tell me how 2+3 becomes =8

OpenStudy (anonymous):

its a puzzle

OpenStudy (anonymous):

i guess its 2^3=8

OpenStudy (anonymous):

48 x 7 = 336

OpenStudy (anonymous):

8 = 4 x 2 27 = 9 x 3 32 = 8 x 4 60 = 12 x 5 72 = 12 x 6 336 = 48 x 7

OpenStudy (anonymous):

then wht about 3+7=27 @nitz

OpenStudy (anonymous):

sorry it should be 14x 7 = 98

OpenStudy (anonymous):

98 is the answer

OpenStudy (anonymous):

bt what is the logic @abb50

OpenStudy (anonymous):

(n*b)+n*x=c b=n+1 x=n-1 n*n+1+n*n-1=c 7+8=c? 7*(7+1)+7*(7-1)=c =56+42=c c=98 7+8=98

OpenStudy (anonymous):

8 = (3 + 1) x 2 27 = (7+2) x 3 32 = (5+3) x 4 60 = (8 + 4)x5 72 = (7 +5)x6 89 = (8 +6) x7

OpenStudy (anonymous):

did i pass the audition @KRAZZy ?

OpenStudy (anonymous):

(2*3)+(1*2)=8 (3*7)+(2*3)=27 (4*5)+(3*4)=32 Similarly (7*8)+(6*7)=98

OpenStudy (anonymous):

not 89 its 98 bt u r ryt @abb50

OpenStudy (anonymous):

sorry. that was a typo.

OpenStudy (anonymous):

nitz i cant understand hw u do this @nitz

OpenStudy (anonymous):

ANOTHER WAY: (2+3)-1 = 4*(2)=8 (3+7)-1 = 9*(3)=27 so (7+8)-1 =14*(7)=98 answer :)

OpenStudy (anonymous):

yeah its better thnx @nitz